the length of a rectangle is 5 inches more than the width. the perimeter of the rectangle is 50 inches. what aer the length and width of the rectangle?
Answer by avani(25)

the length of a rectangle is 5 inches more than the width. the perimeter of the rectangle is 50 inches. what aer the length and width of the rectangle?
since the length is 5" more than width, Let length(L)=x+5 inches
Let, widht(W)=5 inches
perimeter of a rectangle = 2(L) + 2(W)
50 = 2(x+5) + 2(5)
50 = 2x + 10 + 10
50 - 20 = 2x
30/2 = x
15 = x
Thus, length(L) = 15+5 = 20 inches and width = 5 inches
Check:
perimeter = 2(L) + 2(W)= 2(20) + 2(5) = 50 inches